mike is now 5 years older than chelle. in three years, the sum of their ages will 43 how old is mike now?
Answer by TimothyLamb(4379)

x = mike age now
y = chelle age now
---
x = y + 5
x + 3 + y + 3 = 43
---
put the system of linear equations into standard form
---
x - y = 5
x + y = 37
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = mike age now = 21
y = chelle age now = 16
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php